# APCodec: A Neural Audio Codec with Parallel Amplitude and Phase Spectrum Encoding and Decoding

### v1

@@ IV-D Primary Experimental Results
Specifically, we first compared the HiFi-Codec and APCodec because both of them were non-streamable. As shown in Table I, at a sampling rate of 48 kHz and a bitrate of 6 kbps, the proposed APCodec outperformed the HiFi-Codec significantly for all metrics. From the perspective of LSD and MCD, the APCodec exhibited higher quality in decoded audio amplitude spectrum, highlighting its advantage in explicitly modeling amplitude spectra. Similarly, according to the results of the AWPD metrics, it is evident that explicit modeling of phase spectra in the APCodec contributed to improving the precision of decoded phases. However, among the three specific metrics of AWPD, the difference reflected by $\text{AWPD}_{\text{IP}}$ was more pronounced. Although the APCodec outperformed the HiFi-Codec in both $\text{AWPD}_{\text{GD}}$ and $\text{AWPD}_{\text{IAF}}$ metrics, the numerical difference was very small. Overall, the $\text{AWPD}_{\text{GD}}$ values for all the codecs were concentrated in the range of 1.40 to 1.44, while the $\text{AWPD}_{\text{IAF}}$ values were concentrated in the range of 1.44 to 1.47. Additionally, we also observed that, apart from the proposed APCodec and APCodec-S, the $\text{AWPD}_{\text{IP}}$ values for other baseline codecs were all around 1.80. The shared characteristic among these codecs is the absence of explicit phase modeling. Consequently, we hypothesize that their $\text{AWPD}_{\text{IP}}$ values may reflect a consistent initial phase error. The APCodec we proposed achieved a reduction of approximately 0.12 in the $\text{AWPD}_{\text{IP}}$ value through explicit prediction and optimization of the phase. The aforementioned findings suggested a similarity in the phase spectrum continuity of the decoded audio across these codecs according to the results of $\text{AWPD}_{\text{GD}}$ and $\text{AWPD}_{\text{IAF}}$. Nevertheless, the APCodec standed out by producing audio with instantaneous phase values that closely aligned with the natural phase, showcasing superior quality in decoded phase. The aforementioned advantages also contributed to APCodec’s leading position in intelligibility (i.e., STOI) and overall audio quality (i.e., ViSQOL). In terms of generation efficiency, as shown in Table II, whether on GPU or CPU, the APCodec exhibited faster generation speed than the HiFi-Codec. This advantage was more pronounced when running on CPU. This phenomenon indicated that without the parallel acceleration of GPU, utilizing spectra as encoding and decoding objects can significantly enhance generation efficiency compared to the direct encoding and decoding of waveforms. Especially for high-sampling-rate audio codecs, the spectrum-based approach was more suitable due to the larger number of waveform samples. Furthermore, the APCodec was a lightweight model, with a model size of only 26.9% compared to that of the the HiFi-Codec. Therefore, despite the HiFi-Codec [29] incorporating GRVQ to reduce bitrates, its performance in low bitrate scenarios was not as effective as our proposed APCodec.
Then, we compared the APCodec-S with other streamable codecs, i.e., SoundStream, Encodec and AudioDec at 48 kHz and 6 kbps. To ensure the generation of at least one frame of latent code, the minimum length of the input audio for these codecs is 320 samples. Therefore, the latency for these codecs is approximately 6.67 ms for 48 kHz audio. As shown in Table I, the proposed APCodec-S significantly outperformed the baseline SoundStream and Encodec on all metrics. However, we found that the AudioDec, which is a combination of encoder and vocoder, served as a robust baseline, with a ViSQOL score 0.05 higher than the APCodec-S. Yet, it lagged behind the APCodec-S in all other metrics. As illustrated in Table II, the generation speed of the SoundStream and Encodec was relatively fast, slightly trailing behind the APCodec-S, but their model size was 1.78 times that of the APCodec-S. The generation speed of the AudioDec on CPU was relatively slow, just reaching the real-time standard. This may be attributed to the introduction of the HiFi-GAN vocoder. The introduction of the vocoder also resulted in a large model size, approximately 6.4 times that of the APCodec-S. Furthermore, the two-stage training paradigm of the AudioDec also led to operational complexity, in contrast to our proposed end-to-end APCodec.
By comparing the APCodec and APCodec-S at 48 kHz and 6 kbps in Table I, the overall performance of the streamable model decreased compared to non-streamable one. This is reasonable, as the streamable model did not leverage future information. The APCodec can be considered as an upper-bound model for the APCodec-S. Nevertheless, the APCodec-S still outperformed numerous streamable baselines. It’s worth mentioning that the APCodec-S had an increase of 0.06 in $\text{AWPD}_{\text{IP}}$ metric compared to the APCodec. Although this difference was small, during the training process, we clearly observed that the AW-IP loss reduction was challenging for the APCodec-S. This also reflects that in our proposed model, the convergence status of the AW-IP loss can be used to preliminarily estimate the quality of the decoded audio, which is helpful for model selection. Finally, compared to the APCodec, the APCodec-S showed improved efficiency and a further reduction in model size, as shown in Table II. This is because, in the process of transforming the non-streamable model into a streamable one, we chose to replace non-causal convolutions with feed-forward layers instead of causal convolutions used in SoundStream, Encodec and AudioDec. This reduced the model complexity and further improved the generation speed.
@@
TABLE II: Experimental results of generation speed and model complexity for compared codecs on the test set of the VCTK dataset at 48 kHz sampling rate and two bitrates. Here, “$a\times$” represents $a\times$ real time. The bold and underline numbers indicate the optimal and sub-optimal results, respectively.
Codec Bitrate RTF (GPU)$\downarrow$ RTF (CPU)$\downarrow$ Model $\downarrow$
SoundStream 6 kbps 0.0148 (67.6$\times$) 0.231 (4.33$\times$) 83.2M
Encodec 0.0149 (67.1$\times$) 0.232 (4.31$\times$) 83.2M
HiFi-Codec 0.0178 (56.2$\times$) 0.878 (1.14$\times$) 243M
AudioDec 0.0135 (74.1$\times$) 0.771 (1.30$\times$) 108M
APCodec 0.0112 (89.3$\times$) 0.173 (5.78$\times$) 65.4M
APCodec-S 0.0109 (91.7$\times$) 0.112 (8.93$\times$) 46.8M
SoundStream 12 kbps 0.0160 (62.5$\times$) 0.236 (4.24$\times$) 99.2M
Encodec 0.0157 (63.7$\times$) 0.238 (4.20$\times$) 99.2M
HiFi-Codec 0.0178 (56.2$\times$) 0.921 (1.09$\times$) 243M
AudioDec 0.0132 (75.8$\times$) 0.780 (1.28$\times$) 110M
APCodec 0.0120 (83.3$\times$) 0.181 (5.52$\times$) 66.0M
APCodec-S 0.0119 (84.0$\times$) 0.116 (8.62$\times$) 47.4M
@@
To further assess the performance of the codecs at different bitrates, we conducted experiments on these comparative codecs at 48 kHz and 12 kbps. The results are also presented in Tables I and II. For the same codec, there was a noticeable improvement in aspects such as amplitude spectrum quality, intelligibility, overall audio quality and phase spectrum quality at 12 kbps compared to those at 6 kbps. Simultaneously, the former experienced a decrease in generation speed, coupled with an increase in model complexity. This is reasonable, as increasing the number of VQs can reduce the quantization error and increase trainable parameters. The comparison results for different codecs at 12 kbps were essentially consistent with those at 6 kbps. Notably, the ViSQOL score for the APCodec at 12 kbps reached an impressive 4.26 (the maximum score is 4.75). Slightly differently, the performance of the HiFi-Codec at 12 kbps was significantly stronger than that at 6 kbps. Specifically, at 12 kbps, the HiFi-Codec even achieved the highest intelligibility, according to the results of STOI. In addition, the AudioDec also demonstrated a noticeable performance improvement at 12 kbps. This result aligns with expectations, as the AudioDec [33] was originally designed as a 48 kHz audio codec for 12.8 kbps. This observation underscored the suitability of the proposed APCodec for encoding and decoding at low bitrates, showcasing enhanced audio compression capabilities.
@@
Due to their original design as low sampling rate (e.g., 16 kHz or 24 kHz) audio codecs of the SoundStream, Encodec and HiFi-Codec, we also conducted comparative experiments at sampling rates of 16 kHz and 24 kHz. For simplicity, the streamable APCodec was not included in the experiments, and the AudioDec was also excluded due to its original design for 48 kHz. The experimental results are shown in Table I, where the ViSQOL at 24 kHz is not calculated as it is not supported at that sampling rate. It can be observed that both the SoundStream and Encodec still exhibited significant gaps compared to the APCodec at these two sampling rates. At 16 kHz, the APCodec surpassed the HiFi-Codec across all metrics. However, at 24 kHz, the APCodec did not perform as well as the HiFi-Codec in terms of MCD and STOI. Furthermore, it can be noted that in some other metrics (e.g., ViSQOL score), the difference between the HiFi-Codec and APCodec was not as pronounced as it was at 48 kHz. This may be attributed to the fact that the HiFi-Codec is originally designed for 16 kHz and 24 kHz codecs. The above results indicated that while our proposed APCodec exhibited a more pronounced advantage at 48 kHz, applying it at lower sampling rates also yielded good performance.
@@
Based on the above experimental results, we can conclude that APCodec, by leveraging the advantages of parametric codec and waveform codec, is better suited for audio encoding and decoding at both high sampling rates and low bitrates. The APCodec possesses the advantages of high decoded audio quality, low compression rate, fast generation speed, low model complexity, and low latency.
IV-E Analysis and Discussion
@@ IV-E1 Ablation Studies
TABLE III: Experimental results of amplitude spectrum quality, intelligibility, overall audio quality and phase spectrum quality for ablated codecs on the test sets of the VCTK dataset for sampling rate of 48 kHz and bitrate of 6 kbps. The bold numbers indicate the optimal results.
@@
The results of the ablation experiments are shown in Table III. For simplicity, only the LSD, STOI, ViSQOL and $\text{AWPD}_{\text{IP}}$ metrics were used. By comparing the APCodec and APCodec w/o CNV, it can be observed that replacing the modified ConvNeXt v2 network with RCNet resulted in a significant decrease in all metrics. Specifically, the APCodec w/o CNV’s ViSQOL score decreased by 0.5 compared to the APCodec, indicating a significant distortion in the overall audio quality. Additionally, it’s worth noting that the APCodec w/o CNV’s $\text{AWPD}_{\text{IP}}$ metric became as poor as the SoundStream, Encodec, HiFi-Codec and AudioDec. This indicates that the RCNet impeded the learning of phase information, contradicting the conclusions in [37] and [38]. Hence, the RCNet was apt for tasks involving phase predictions [37] and vocoders [28, 38], leveraging its cumulative dilated convolutional layers to broaden the receptive field. Nevertheless, in codec tasks that necessitated a more sophisticated phase modeling, the RCNet exhibited an inadequate modeling capability. The ConvNeXt v2 network, borrowed from the field of image processing, exhibited stronger modeling capabilities, making it well-suited for the design of codec models.
@@
Regarding the ablation studies on some training strategies, by comparing the APCodec and APCodec w/o MelMSE, it is evident that the MSE loss on the mel spectrogram had a positive impact on intelligibility and overall audio quality. The MSE exhibited greater sensitivity to outliers when compared to MAE. It can be viewed as a complement to MAE, collectively enhancing the overall quality of the mel spectrogram. Derived from the outcomes of the LSD and $\text{AWPD}_{\text{IP}}$, it is evident that the influence of this MSE loss on amplitude and phase learning was contradictory. Enhancing the quality of the amplitude spectrum unavoidably came at the expense of compromising the quality of the phase spectrum. By comparing the APCodec and APCodec w/o QLoss, it can be observed that the quantization loss had a significant impact on the amplitude spectrum quality, intelligibility, and overall audio quality, while its influence on the phase spectrum quality was relatively minor. The incorporation of quantization loss effectively alleviated quantization errors, thereby contributing to the enhancement of APCodec’s performance. Replacing MRD with MSD significantly impacted the amplitude spectrum quality and overall audio quality, based on the results of the APCodec w/o MRD. This is in line with expectations because the MRD focused more on the quality of the amplitude spectrum, making it suitable for our spectrum-based approach. Finally, the hinge-form adversarial loss was more effective compared to the least-squares-form adversarial loss commonly used in some vocoder tasks [28, 38], according to the results of the APCodec w/o Hinge. In terms of auditory sensation, the APCodec w/o Hinge experienced very apparent harsh noise.
@@
In terms of the role of the proposed knowledge distillation strategy for the streamable APCodec, we compared the APCodec-S and APCodec-S w/o KD. The results are also listed in Table III. It can be observed that without the guidance of the non-streamable teacher model, the streamable student model exhibited slight decreases across all metrics. In particular, the $\text{AWPD}_{\text{IP}}$ of the APCodec-S w/o KD deteriorated to the level of the initial phase error, resembling the patterns seen in SoundStream, Encodec, HiFi-Codec and AudioDec. This indicates that the low-latency implementation on model structures discussed in Section III-C hindered the learning of phase, and accurate phase prediction required a network with a broader receptive field. The knowledge distillation strategy can effectively alleviate the challenge of phase learning difficulty ($\text{AWPD}_{\text{IP}}$ reduced by 0.05), thereby promoting overall audio quality improvement.
@@
TABLE IV: Experimental results of amplitude spectrum quality, overall audio quality and phase spectrum quality for the comparison between HiFi-Codec and APCodec and the comparison between AudioDec and APCodec-S on the test sets of the Opencpop dataset and FSD50K dataset for sampling rate of 48 kHz and bitrate of 6 kbps. The bold numbers indicate the optimal results.
@@
Opencpop HiFi-Codec 1.08 3.52 1.80
APCodec 0.864 4.21 1.67
AudioDec 0.967 4.10 1.81
@@
FSD50K HiFi-Codec 1.17 3.18 1.81
APCodec 0.845 3.93 1.71
AudioDec 1.01 3.75 1.82
APCodec-S 0.874 3.84 1.76
IV-E2 Validation on Diverse Audio Datasets

### v2

@@ III-D Primary Experimental Results
Specifically, we first compared the proposed APCodec with DAC because both of them were non-streamable. As shown in Table I, at a sampling rate of 48 kHz and a bitrate of 6 kbps, the proposed APCodec outperformed the DAC significantly for most metrics, except for the STOI metric. From the perspective of LSD and MCD, the APCodec exhibited higher quality in decoded audio amplitude spectrum, highlighting its advantage in explicitly modeling amplitude spectra. Similarly, according to the results of the AWPD metrics, it is evident that explicit modeling of phase spectra in the APCodec contributed to improving the precision of decoded phases. However, among the three specific metrics of AWPD, the difference reflected by $\text{AWPD}_{\text{IP}}$ was more pronounced. The $\text{AWPD}_{\text{GD}}$ values for all codecs at 48 kHz and 6 kbps in Table I were concentrated in the range of 1.40 to 1.44, while the $\text{AWPD}_{\text{IAF}}$ values were concentrated in the range of 1.44 to 1.47. The differences in $\text{AWPD}_{\text{GD}}$ and $\text{AWPD}_{\text{IAF}}$ metrics between different codecs were minor. Additionally, we also observed that, apart from the proposed APCodec and APCodec-S, the $\text{AWPD}_{\text{IP}}$ values for other baseline codecs were all around 1.80. The shared characteristic among these codecs is the absence of explicit phase modeling. Consequently, we hypothesize that their $\text{AWPD}_{\text{IP}}$ values may reflect a consistent initial phase error. The APCodec we proposed achieved a reduction of approximately 0.12 in the $\text{AWPD}_{\text{IP}}$ value through explicit prediction and optimization of the phase. The aforementioned findings suggested a similarity in the phase spectrum continuity of the decoded audio across these codecs according to the results of $\text{AWPD}_{\text{GD}}$ and $\text{AWPD}_{\text{IAF}}$. Nevertheless, the APCodec standed out by producing audio with instantaneous phase values that closely aligned with the natural phase, showcasing superior quality in decoded phase. Although the APCodec lagged behind the DAC in intelligibility (i.e., STOI), it benefited from the aforementioned advantages, placing it in a leading position in terms of overall audio objective quality (i.e., ViSQOL). In terms of generation efficiency, as shown in Table II, whether on GPU or CPU, the APCodec exhibited faster generation speed than the DAC. This advantage was more pronounced when running on CPU. The generation speed of APCodec on the CPU was approximately 14 times faster than that of the DAC, and the DAC was unable to achieve real-time generation on CPU. This phenomenon indicated that without the parallel acceleration of GPU, utilizing spectra as coding objects can significantly enhance generation efficiency compared to the direct encoding and decoding of waveforms. Especially for high-sampling-rate audio codecs, the spectrum-based approach was more suitable due to the larger number of waveform samples. Furthermore, the APCodec was a lightweight model, with a model size of only 23.2% compared to that of the the DAC. Although the APCodec and DAC had similar subjective perceptual quality, according to the MUSHRA scores in Table II, overall, the APCodec performed the best. This is because it demonstrated significantly faster generation speed, a lighter model, and superior performance on most objective metrics.
@@
TABLE II: Objective and subjective experimental results for compared codecs on the test set of the VCTK dataset at 48 kHz sampling rate and two bitrates. Here, “$a\times$” represents $a\times$ real time. The bold and underline numbers indicate the optimal and sub-optimal results, respectively.
Codec Bitrate RTF (GPU)$\downarrow$ RTF (CPU)$\downarrow$ Model $\downarrow$ MUSHRA score
APCodec 6 kbps 0.0112 (89.3$\times$) 0.173 (5.78$\times$) 65.4M 88.48$\pm$0.87
APCodec-S 0.0109 (91.7$\times$) 0.112 (8.93$\times$) 46.8M 88.07$\pm$0.89
Encodec [26] 0.0149 (67.1$\times$) 0.232 (4.31$\times$) 83.2M 85.91$\pm$1.25
AudioDec [33] 0.0132 (75.8$\times$) 0.771 (1.30$\times$) 108M 88.34$\pm$0.89
DAC [34] 0.0195 (51.3$\times$) 2.47 (0.405$\times$) 282M 88.28$\pm$0.92
APCodec 12 kbps 0.0120 (83.3$\times$) 0.181 (5.52$\times$) 66.0M 90.68$\pm$0.86
APCodec-S 0.0119 (84.0$\times$) 0.116 (8.62$\times$) 47.4M 90.16$\pm$0.85
Encodec [26] 0.0157 (63.7$\times$) 0.238 (4.20$\times$) 99.2M 88.09$\pm$1.21
AudioDec [33] 0.0135 (74.1$\times$) 0.780 (1.28$\times$) 110M 89.66$\pm$0.93
DAC [34] 0.0216 (46.3$\times$) 2.68 (0.373$\times$) 283M 90.78$\pm$0.91
Then, we compared the APCodec-S with other streamable codecs, i.e., Encodec and AudioDec at 48 kHz and 6 kbps. As mentioned in Section II-C, these codecs all had an unavoidable fixed latency. According to the current experimental setup, the latency for APCodec-S, Encodec and AudioDec was all approximately 6.67 ms (i.e., 320 samples) for 48 kHz audio, hence their comparison is fair. As shown in Table I and II, the proposed APCodec-S significantly outperformed the baseline Encodec on all objective and subjective metrics. However, we found that the AudioDec, which is a combination of encoder and vocoder, served as a robust baseline, with an objective ViSQOL score 0.05 higher than the APCodec-S and a similar subjective MUSHRA score with the APCodec-S. Yet, it lagged behind the APCodec-S in all other metrics. As illustrated in Table II, the generation speed of the Encodec was relatively fast, slightly trailing behind the APCodec-S, but its model size was 1.78 times that of the APCodec-S. The generation speed of the AudioDec on CPU was relatively slow, just reaching the real-time standard. This may be attributed to the introduction of the HiFi-GAN vocoder. The introduction of the vocoder also resulted in a large model size, approximately 2.3 times that of the APCodec-S. Furthermore, the two-stage training paradigm of the AudioDec also led to operational complexity, in contrast to our proposed end-to-end APCodec.
By comparing the APCodec and APCodec-S at 48 kHz and 6 kbps in Table I, the overall objective performance of the streamable model decreased compared to non-streamable one. This is reasonable, as the streamable model did not leverage future information. The APCodec can be considered as an upper-bound model for the APCodec-S. Nevertheless, the APCodec-S still outperformed numerous streamable baselines. It’s worth mentioning that the APCodec-S had an increase of 0.06 in $\text{AWPD}_{\text{IP}}$ metric compared to the APCodec. Although this difference was small, during the training process, we clearly observed that the AW-IP loss reduction was challenging for the APCodec-S. This also reflects that in our proposed model, the convergence status of the AW-IP loss can be used to preliminarily estimate the quality of the decoded audio, which is helpful for model selection. Although the introduction of low-latency implementation led to a significant deterioration in objective metrics, according to the MUSHRA scores in Table I, the subjective perceptual quality only slightly declined. Furthermore, compared to the APCodec, the APCodec-S showed improved efficiency and a further reduction in model size, as shown in Table II. This is because, in the process of transforming the non-streamable model into a streamable one, we chose to replace non-causal convolutions with feed-forward layers instead of causal convolutions used in Encodec and AudioDec. This reduced the model complexity and further improved the generation speed.
@@
To further assess the performance of the codecs at different bitrates, we conducted experiments on these comparative codecs at 48 kHz and 12 kbps. The results are also presented in Tables I and II. For the same codec, there was a noticeable improvement in both objective and subjective aspects at 12 kbps compared to those at 6 kbps. Simultaneously, the former experienced a decrease in generation speed, coupled with an increase in model complexity. This is reasonable, as increasing the number of VQs can reduce the quantization error and increase trainable parameters. The comparison results for different codecs at 12 kbps were essentially consistent with those at 6 kbps. Notably, the ViSQOL score for the APCodec at 12 kbps reached an impressive 4.26 (the maximum score is 4.75). However, the performance of other codecs has also significantly improved. For instance, the DAC achieved remarkably high intelligibility for audio decoded at 12 kbps, as indicated by the STOI results, despite it was inferior to or comparable with our proposed APCodec in other metrics. In addition, the AudioDec also demonstrated a noticeable performance improvement at 12 kbps. This result aligns with expectations, as the AudioDec [33] was originally designed as a 48 kHz audio codec for 12.8 kbps. Fortunately, the proposed APCodec-S still maintained its overall superiority over AudioDec at 12 kbps, with the difference in ViSQOL metrics being only 0.01. However, apart from phase metrics, the difference between APCodec and other codecs at 12 kbps was smaller compared to the difference at 6 kbps across other metrics. This observation underscored the suitability of the proposed APCodec for encoding and decoding at low bitrates, showcasing enhanced audio compression capabilities.
@@
Due to their original design as low sampling rate for some well-known audio codecs, e.g., SoundStream, Encodec and HiFi-Codec, we also conducted comparative experiments at sampling rates of 16 kHz and 24 kHz. The objective experimental results are shown in Table I. It can be observed that both the streamable Encodec and SoundStream exhibited significant gaps compared to our proposed streamale APCodec-S at these two sampling rates. For comparisons between non-streamable codecs, at 16 kHz, the APCodec surpassed the HiFi-Codec across all metrics. However, at 24 kHz, the APCodec did not perform as well as the HiFi-Codec in terms of MCD and STOI. This may be attributed to the fact that the HiFi-Codec originally excelled at a 24 kHz sampling rate [29]. Interestingly, when comparing the APCodec and APCodec-S at low sampling rates and high bitrates, the APCodec-S even outperformed APCodec in terms of MCD and ViSQOL metrics, which suggested an improvement in perceptual quality in the mel scale. This means that the proposed low-latency implementation is more effective for low-sampling-rate APCodec, because the low-latency implementation under high sampling rate conditions obviously reduced the ViSQOL score as shown in Table I. The above results indicated that while our proposed APCodec exhibited a more pronounced advantage at 48 kHz, applying it at lower sampling rates also yielded good performance.
@@
Based on the above experimental results, we can conclude that APCodec, by leveraging the advantages of parametric codec and waveform codec, is better suited for audio coding at both high sampling rates and low bitrates. The APCodec possesses the advantages of high decoded audio quality, high compression rate, fast generation speed, low model complexity, and low latency.
III-E Analysis and Discussion
@@ III-E1 Ablation Studies
TABLE III: Objective experimental results for ablated codecs on the test sets of the VCTK dataset for sampling rate of 48 kHz and bitrate of 6 kbps. The bold numbers indicate the optimal results.
@@
The results of the ablation experiments are shown in Table III. For simplicity, only the LSD, STOI, ViSQOL and $\text{AWPD}_{\text{IP}}$ metrics were used. By comparing the APCodec and APCodec w/o CNV, it can be observed that replacing the modified ConvNeXt v2 network with RCNet resulted in a significant decrease in all metrics. The APCodec w/o CNV’s ViSQOL score decreased by 0.5 compared to the APCodec, indicating a significant distortion in the overall audio quality. Specifically, according to the results of LSD and $\text{AWPD}_{\text{IP}}$, the RCNet impeded the learning of both amplitude and phase spectra, contradicting the conclusions in [39]. Hence, the RCNet was apt for tasks involving vocoders [28, 39], leveraging its cumulative dilated convolutional layers to broaden the receptive field. Nevertheless, in codec tasks that necessitated a more sophisticated parallel amplitude and phase modeling, the RCNet exhibited an inadequate modeling capability. The ConvNeXt v2 network, borrowed from the field of image processing, exhibited stronger modeling capabilities, making it well-suited for the design of codec models.
@@
Regarding the ablation studies on some training strategies, by comparing the APCodec and APCodec w/o MelMSE, it is evident that the MSE loss on the mel spectrogram had a positive impact on intelligibility and overall audio quality. The MSE exhibited greater sensitivity to outliers when compared to MAE. It can be viewed as a complement to MAE, collectively enhancing the overall quality of the mel spectrogram. Derived from the outcomes of the LSD, it is reasonable that removing the amplitude-related mel-spectrogram MSE loss would lead to worse LSD. However, metric $\text{AWPD}_{\text{IP}}$ has improved, which might be because removing the mel-spectrogram MSE loss increased the weight of the phase loss. By comparing the APCodec and APCodec w/o QLoss, it can be observed that the quantization loss had a significant impact on the amplitude spectrum quality, intelligibility, and overall audio quality, while its influence on the phase spectrum quality was relatively minor. The incorporation of quantization loss effectively alleviated quantization errors, thereby contributing to the enhancement of APCodec’s performance. Replacing MRD with MSD significantly impacted the amplitude spectrum quality and overall audio quality, based on the results of the APCodec w/o MRD. This is in line with expectations because the MRD focused more on the quality of the amplitude spectrum, making it suitable for our spectrum-based approach. Finally, the hinge-form adversarial loss was more effective compared to the least-squares-form adversarial loss commonly used in some vocoder tasks [28, 39], according to the results of the APCodec w/o Hinge. Although replacing the adversarial loss form resulted in a 0.004 increase in STOI, this difference was not significant according to a $t$-test, while the ViSQOL value significantly decreased. This indicates that the hinge-form, compared to the least-squares-form form, does not improve intelligibility but significantly enhances audio quality. In terms of auditory sensation, the APCodec w/o Hinge experienced very apparent harsh noise.
@@
In terms of the role of the proposed knowledge distillation strategy for the streamable APCodec, we compared the APCodec-S and APCodec-S w/o KD. The results are also listed in Table III. It can be observed that without the guidance of the non-streamable teacher model, the streamable student model exhibited slight decreases across all metrics. In particular, the $\text{AWPD}_{\text{IP}}$ of the APCodec-S w/o KD deteriorated to the level of the initial phase error, resembling the patterns seen in Encodec, AudioDec and DAC. This indicates that the low-latency implementation on model structures discussed in Section II-C hindered the learning of phase, and accurate phase prediction required a network with a broader receptive field. The knowledge distillation strategy can effectively alleviate the challenge of phase learning difficulty ($\text{AWPD}_{\text{IP}}$ reduced by 0.05), thereby promoting overall audio quality improvement.
@@
TABLE IV: Objective experimental results for the comparison between APCodec and DAC and the comparison between APCodec-S and AudioDec on the test sets of the Common Voice dataset, Opencpop dataset and FSD50K dataset for sampling rate of 48 kHz and bitrate of 6 kbps, when finetuning on these three datasets individually. The bold numbers indicate the optimal results.
@@
Common Voice APCodec 0.872 4.25 1.73
DAC [34] 0.955 4.08 1.78
APCodec-S 0.838 4.08 1.75
AudioDec [33] 0.929 4.10 1.80
Opencpop APCodec 0.864 4.21 1.67
DAC [34] 0.972 3.94 1.77
@@
AudioDec [33] 0.967 4.10 1.81
FSD50K APCodec 0.853 3.95 1.71
DAC [34] 0.929 3.86 1.78
APCodec-S 0.887 3.86 1.76
AudioDec [33] 1.04 3.77 1.82
III-E2 Validation on Diverse Audio Datasets